# RRM: Robust Reward Model Training Mitigates Reward Hacking

### v1

@@ Appendix A Appendix
A.1 Proofs
Proof of Proposition 3.1
Proof.
We show the conclusion with a special case. Assume $S=s(X,Y_{1},Y_{2})+\epsilon_{s}$ with $\epsilon_{s}\sim N(0,\sigma_{s})$, and $A=a(X,Y_{1},Y_{2})+\epsilon_{a}$ with $\epsilon_{a}\sim N(0,\sigma_{a})$. $\mathbb{P}(C=1)=\sigma(\beta_{s}S+\beta_{a}A+\alpha+\epsilon_{c})$ with constants $\beta_{s},\beta_{a},\alpha\in\mathbb{R}$ and random error $\epsilon_{c}\perp(S,A)$. If $\beta_{a}=0$, then $\mathcal{H}_{0}$ is true. If $\beta_{a}=1$, then $\mathcal{H}_{1}$ is true. In extreme case that $Corr(S,A)=1$, then $A=\beta_{as}S+\alpha_{a}$ for some constants $\alpha_{a}\in\mathbb{R}$ and $\beta_{as}\in\mathbb{R}^{+}$. Then the model cannot tell if $\beta_{a}=0$ or not. This is because when $\beta_{a}=0$, we can still reparametrize it as $\mathbb{P}(C=1)=\sigma((\beta_{s}-\beta_{as})S+A+(\alpha-\alpha_{a})+\epsilon_{c})$. ∎
Proof of Proposition 3.2
Proof.
A.2 Reward Model Training Datasets
We reuse the datasets in RLHFlow (Dong et al., 2024). We list the data sources and number of examples in Table 3. The authors of the original paper delete the samples with similar scores when the scores are available because when the model is well calibrated, these samples are more likely to mislabelled. Thus the total number is smaller than the sum of each individual datasets. We show the length distribution of chosen and rejected responses from each individual data source in Figure 6. HH-RLHF, SHP, HelpSteer, UltraFeedback show bias towards rejected responses as the first length bin. SHP, HelpSteer, and UltraFeedback have longer chosen responses than rejected ones on average.
Source Number of Examples
Table 3: Composition of reward model training dataset
@@ A.2 Reward Model Training Datasets
A.3 Data augmentation python code
@@
A.4 Training details for ODIN
We use the same loss as described in Chen et al. . We train Gemma-2-9b-it for 1 epoch on the same data we used for RM. AdamW is our optimizer and the learning rate is set to 2e-6 with cosine scheduler. We use Flash-Attention to accelerate the training while applying the Deepspeed Zero-Stage 3 to get batch_size=16 on each GPU (the global batch size is 128) to make sure the calculation of the Pearson correlation between the head value and the length of the responses is stable.
A.5 Preliminaries in Causal inference

### v2

@@ Appendix A Appendix
A.1 Additional length analysis of reward model training datasets
In this section, we show the length distribution of chosen and rejected responses from each individual data source in Figure 6. HH-RLHF, SHP, HelpSteer, UltraFeedback show bias towards rejected responses as the first length bin. SHP, HelpSteer, and UltraFeedback have longer chosen responses than rejected ones on average.
@@ A.1 Additional length analysis of reward model training datasets
A.2 Data augmentation python code
@@
A.3 Training details for ODIN
We use the same loss as described in Chen et al. . We train Gemma-2-9b-it for 1 epoch on the same data we used for RM. AdamW is our optimizer and the learning rate is set to 2e-6 with cosine scheduler. We use Flash-Attention to accelerate the training while applying the Deepspeed Zero-Stage 3 to get batch size 16 on each GPU (the global batch size is 128) to make sure the calculation of the Pearson correlation between the head value and the length of the responses is stable.
A.4 Preliminaries in causal inference
@@
A.5 Additional Results with Gemma-2-2b-it
To further verify effectiveness of our approach, we train Gemma-2-2b-it reward model (RM) and robust reward model (RRM), respectively. Table 4 shows the results on the reward bench. RRM again shows improvement on the Chat Hard. It shows some regression on Safety and Reasoning, where we hypothesis that some context-free nature of reasoning and safety makes the RRM perform worse. Overall, RRM shows positive effect. We also test the reward models on AlpacaEval2 using best-of-n policies. The results are shown in Table 5. We use the trained reward model to rank the responses generated from the Gemma-2-9b-it model, and observe consistent gains on RRM over RM.
Model Chat Chat Hard Safety Reasoning Average
RM 96.49 42.54 72.70 72.30 71.01
RRM 97.21 49.01 72.71 70.08 72.25
Table 4: Comparison of test accuracy of Reward-Bench. RRM shows improvement upon RM on Chat and Chat Hard with an average 1.75% improvement of accuracy.
Reward Policy AlpacaEval-2
LC (%) ($\uparrow$) WR (%) ($\uparrow$) Length ($\downarrow$)
RM BoN (N=8) 44.21 45.83 2264
RRM BoN (N=8) 54.37 47.19 1749
RM BoN (N=64) 47.32 52.23 2316
RRM BoN (N=64) 60.42 56.50 1870
Table 5: Comparison among different reward models on various aligned policies. WR stands for win-rate against GPT-4. LC stands for length-controlled win-rate. Length is the average number of characters in the generated responses. RRM shows quality improvements with shorter responses over RM.
A.6 Additional Analysis with Mixed Artifacts
To investigate the effect of RRM on mixed artifacts, we conduct an additional experiment as follows:
  1. 1.
    With $p=0.1$, wrap the whole chosen response with “**” as bold-face.
  2. 2.
[…]
  3. 3.
    Train RM and RRM on the above dataset.
[…]
(a) Best of 8 responses
(b) Best of 64 responses
Figure 7: Proportion of BoN generated responses with emoji versus the rate of injected emoji. For each policy, we first sample $N$ ($N=8$ or $64$) responses on AlpacaEval-2 prompts, then append emoji after each response with probability (Rate) 5%, 10%, 20%, 50%, respectively. Then we compute the proportion of BoN responses that have the above artifact (Artifact). The BoN policies induced by RRM are more robust to artifacts injected in the responses, suggesting that the proposed approach enables the model to focus more on the contextual signals instead of context-free artifacts in the reward model training data.
A.7 Discussion on Data Filtering Strategies
In this work, we assume the preference labels should be purely controlled by the prompt dependent signal. However, there can be cases such that prompt-independent signals can contribute to the preference label. For example, a responsible AI should be always safe and cares about the diversity. In the data augmentation, we sometimes use rejected response as new chosen in “Non-contextuals”. See the last four triplets in “Non-contextuals” of Equation 5. For “Neutrals”, we also assume 0.5 winning probability of a non-contextual response pair. These treatment may cause the unwanted behavior of AI if we use unsafe response as the new chosen or assigning winning probability of 0.5 on a pair of (safe, unsafe) responses.
To address this, we have a few treatments that can be applied in future works:
  • Use a trained safety pointwise or Bradley-Terry model to filter out triplets that has low safety scores on the chosen responses.
  • Use AI feedback such as Constitutional AI (Bai et al., 2022c) to ensure the augmented triplets have high quality chosen responses and consistent preference according to certain non-contextual rules. The rules can include safety, style, factuality.
Appendix B Ethics Statement
Our research adheres to the ethical guidelines. Our work aims to mitigate reward hacking in RLHF, contributing to the development of more reliable AI systems that better align with human preferences. Our data augmentation process explicitly addresses and mitigates common forms of bias, thus reducing the potential for harm in practical applications of AI systems. Our model was designed with fairness in mind, particularly in avoiding biases related to response length and format, which can unfairly influence AI decision-making.
Appendix C Reproducibility Statement
All code used for training the reward models (RM and RRM) and for running the experiments described in this paper will be made publicly available upon publication. This includes the implementation of the data augmentation pipeline, the reward model training process, and policy alignment. The datasets used in our experiments are publicly available. We provide the complete set of hyperparameters used for training the models, including learning rates, batch sizes, and other optimization settings. The evaluation approaches are also public available and can be fully reproduced.